What is the Kp of x1- if a 0.0220-M solution of KX has a pH of 11.852? Kb =
Answer #1

00230 M K* Kx + Ft + x = (x ) = 0-0220 m2 - U) pH = 11.85? a por 14-pH = 14-11-852 S a post = 20148 pot = -log[H3 = 2o148 a C